# Locate-then-edit for Multi-hop Factual Recall under Knowledge Editing

### v2

@@ 3.1 How the Pre-edited Model Reasons Fact Recall Tasks
For a multi-hop fact recall task, the knowledge chain is represented as $C=\left(s_{1},r_{1},o_{1}\right)\oplus\dots\oplus\left(s_{n},r_{n},o_{n}\right)$. The model may employ multiple strategies to answer such tasks, including the formation of a single super-relation (Ju et al., 2024) $(s_{1},r_{mul},o_{n})$, where $r_{mul}=r_{1}\to\dots\to r_{n}$, or by segmenting the task into one explicit recall step followed by several implicit recall steps to answer step-by-step. Previous research (Hou et al., 2023) suggests that models typically engage in reasoning by considering each single-hop recall individually.
Based on this understanding, we hypothesize that the model will prioritize deducing the implicit subjects $\{s_{2},\dots,s_{n}\}$ and subsequently recall the final answer $o_{n}$ based on the last implicit subject $s_{n}$. The subsequent sections aim to verify this hypothesis by examining the model’s behavior in structured multi-hop fact recall tasks.
@@
Interpretation via Hidden Representations. We use LogitLens to examine the accumulation of information related to the implicit subject $s_{2}$ and the final answer $o_{2}$ in the two-hop scenario. The model’s predictions for $o_{2}$, are derived from the last token of the prompt, where crucial information about the resolved implicit subject $s_{2}$ should be propagated (Biran et al., 2024). Therefore, we focus on the hidden state $h_{l}$ at the $l$-th layer of the last token position, analyzing $\text{Info}(h_{l},s_{2})$ and $\text{Info}(h_{l},o_{2})$ as measures of the information related to $s_{2}$ and $o_{2}$ contained in $h_{l}$. Intuitively, these metrics quantify how much information about $s_{2}$ and $o_{2}$ accumulates in the hidden state. The results, depicted in Figure 2(a), show that $\text{Info}(h_{l},s_{2})$ gradually reaches its peak during middle layers [15-17], while $\text{Info}(h_{l},o_{2})$ increases and peaks during later layers [21-23]. This pattern suggests that, in multi-hop tasks, the implicit subject $s_{2}$ is processed during the middle layers before reaching the final answer $o_{2}$.
@@
To explore if single-hop fact recalls $(s,r,o)$ follow the same trend as in multi-hop cases, we conducted a similar experiment using LogitLens. The results, shown in Figure 2(b), indicate that $\text{Info}(h_{l},s)$ significantly increases after layer 24 and peaks at layer 27, whereas $\text{Info}(h_{l},o)$ consistently reaches its peak during layers 21,22,23. This finding implies that there is no significant peak for the subject information before the final answer probability begins to accumulate, suggesting that the accumulation process of the final answer in single-hop cases may not be significantly correlated with the subject information at the last token.
@@
(b) Causal Intervention result of MLP hidden state in last token position
@@
Figure 3: Causal Intervention Result: A brighter color signifies a stronger intervention effect. In each subfigure, upper row represents experimental group, while upper row is control group. Note that negative effect values ($\leq$ 0) are clipped to 0 in both groups for better visualization. (a) is probability change $IE_{h}$ of intervention $\mathcal{I}_{h}$, (b) is probability change $IE_{m}$ of intervention $\mathcal{I}_{m}$.
Causal Intervention. Next, we explore whether the appearance of $s_{2}$ guides the subsequent information accumulation process of the final answer $o_{2}$. To this end, we aim to identify which layers facilitate this influence. We propose an intervention experiment where we reduce the information content of $s_{2}$ at the last token position and observe the changes in the output probability of the final answer in the last prediction layer.
Specifically, we replace the hidden state $h_{l}$ in layer $\ell$ of the last token with $h^{*}_{l}$, and the corresponding logits $s_{l}$ ($=W_{U}h_{l}$) and $s^{*}_{l}$ ($=W_{U}h^{*}_{l}$) for $h_{l}$ and $h^{*}_{l}$, respectively. $s^{*}_{l}$ is defined as:
where we minimize the logits corresponding to the tokens in $s_{2}$ without altering the values of other tokens, aiming to diminish the effect of $s_{2}$. This setup allows us to describe the process through a causal intervention framework, where the intervention $\mathcal{I}_{h}$ and the effect $IE_{h}$ are defined as follows:
$\mathcal{I}_{h}:h^{*}_{l}=h_{l}+\operatorname*{arg\,min}_{\Delta h_{l}}\|W_{U}(h_{l}+\Delta h_{l})-s^{*}_{l}\|^{2},\quad IE_{h}=p_{L}[j]-p_{L}^{E}[j],\quad j\in o_{2},$ (2)
where $L$ is the last layer, $p_{L}[j]$ denotes the original output probability of $o_{2}$ in the $L$-th layer, and $p^{E}_{L}[j]$ is the probability after the intervention is applied. This approach illustrates how the hidden states and probabilities are expected to change when the logits are modified to $s^{*}$. For computational efficiency, we opt to approximate $h^{*}_{l}$ using a combination of least squares and minimum-norm methods (Lawson & Hanson, 1995) (further details are provided in Appendix B.2).
For comparison, we also randomly select an irrelevant token $j\notin s_{2}\cup o_{2}$ to execute the intervention as the control group. Figure 3(a) presents the outcomes of our intervention experiments across all layers, where a brighter color signifies a stronger intervention effect. We found a clear positive impact from intervening in layers [17-19] for the experimental group, in contrast to no significant effects observed in the control group across all layers. This suggests that the information of $s_{2}$ encoded in the intermediate layers plays a crucial role in the probability accumulation process of $o_{2}$. We also do the same causal intervention experiments for single-hop fact recall (see Appendix  B.3 for the results). However, the results indicate that the prediction of $o$ does not significantly rely on the subject information $s$ in the single-hop fact recall.
Takeaway 2 Unlike the mechanism of reasoning the knowledge in single-hop scenarios, in the reasoning process of the second-hop knowledge in two-hop scenarios, the accumulated subject information has causal effects on the final answer, guiding the extraction of related knowledge in the last layer.
Intermediate Reasoning Results Influence the Knowledge Extraction from MLP. As previous studies claimed that single-hop tasks retrieve subject information from MLP layers (Meng et al., 2022a; b), we will focus on MLP layers to further investigate the specific mechanisms to answer how the implicit subject $s_{2}$ influences the prediction of the final answer $o_{2}$. We conducted a causal intervention experiment similar to the experiments above but focused specifically on the MLP component. Specifically, we aim to replace $m_{l}$ (the output hidden state of the last token in the $l$-th MLP layer) with $m^{*}_{l}$, where we have $s_{l}=W_{U}m_{l}$ and $s^{*}_{l}=W_{U}m^{*}_{l}$ with $s^{*}_{l}$ is same as in (1). The intervention $\mathcal{I}_{m}$ shares the same idea as in (2), except that $h_{l}$ is replaced with $m_{l}$. However, we redefine the intervention effect $IE_{m}$, which differs from the previous $IE_{h}$. In detail, we no longer use the probability at the final layer as the metric; instead, we use the probability calculated from the output of MLP at the modified layer $l$. In total, our causal intervention is formulated as
$\mathcal{I}_{m}:m^{*}_{l}=m_{l}+\operatorname*{arg\,min}_{\Delta m_{l}}\|W_{U}(m_{l}+\Delta m_{l})-s^{*}_{l}\|^{2},\quad IE_{m}=p_{l}[j]-p_{l}^{E}[j],\quad j\in o_{2}.$
@@
Figure 3(b) presents the outcomes of our intervention experiments across all layers with the similar control group as in the above. [^2] Note that, considering the tiny output probability of MLP, we did not use normalization of probability changes here. The clear positive impact from intervening in the intermediate layers [17-21] is demonstrated in the experimental group, in contrast to negligible effects observed in the control group across all layers. This suggests that the implicit subject $s_{2}$ at the last token position was used for retrieving the related information of $o_{2}$ from later MLP layers. Thus, it plays an important role in the probability accumulation process of $o_{2}$. Note that this is in contrast with previous work (Meng et al., 2022a; 2023), which mentioned that explicit single-hop tasks primarily rely on the subject position token to retrieve information from earlier MLP layers.
Takeaway 3 During the reasoning process of the second-hop knowledge in two-hop scenarios, information related to the subject is used for retrieving relevant knowledge of the final answer from later MLP layers of the last token position, which is from the earlier MLP layers in single-hop cases.
@@ 3.2 Why Existing locate-then-edit KE Methods Failed
Table 1: Comparison of QA and Cloze Formats for $D_{Pre}$ and $D_{Post}$
Edit Batch QA Format(%$\uparrow$ Cloze Format(%$\uparrow$ $D_{Pre}$ $D_{Post}$ $D_{Pre}$ $D_{Post}$ GPT-J $50.62$ $41.72$ $20.31$ $18.63$ Edit=1 64.29 2.93 43.37 $4.60$ Edit=100 $63.27$ $3.35$ $42.86$ 3.35
@@
Based on the findings above, we can provide an explanation for the unsatisfactory performance of the existing locate-then-edit methods. For an editing instance $(s,r,o\to o^{*})$, using only the corresponding explicit single-hop prompt for editing is insufficient as previous methods only update the relevant knowledge in the shallow MLP layers but fail to propagate the changes to deeper layers, which is utilized in multi-hop fact recall tasks.
We provide a concrete example for a better understanding. Given an editing instance $(Spain,captical,Madrid\to Hartford)$, and $Q_{cloze}(s)$ is “The capital city of Spain is”. Existing methods modify the weights of shallow MLPs with $Q_{cloze}(s)$ to make it answer Hartford. The paradigm may be well-suited for cases where the modified information is queried in a single-hop manner, as these tasks retrieve answers from the early MLP layers. However, it will be ineffective when the modified knowledge is queried in the second or later fact recall steps, where the model relies on deeper MLP layers at the last token position for knowledge retrieval. In this example, the first-hop query “The capital city of Spain is located in the continent of” should be answered correctly because it retrieves the knowledge $(Spain,captical,Hartford)$ in shallow MLPs. However, the second one “The capital city of the country has nationals Pablo Picasso is” is still answered with Madrid because the knowledge $(Spain,captical,Madrid)$ stored in later MLPs does not changed.
To verify our above claim, we divide two-hop fact recall tasks into two sets $D_{Pre}$ and $D_{Post}$, depending on the position of the edited knowledge within the two-hop reasoning process. Specifically, for an edited knowledge $\left(s,r,o,o*\right)$, we have the following two sets after editing.
$D_{Pre}$ =$=\{\left(s,r,o^{*}\right)\oplus\left(s_{2},r_{2},o_{2}\right)\},$
$D_{Post}$ =$=\{\left(s_{1},r_{1},o_{1}\right)\oplus\left(s,r,o^{*}\right)\}.$
We sampled two subsets with approximately equal size from the MQuAKE-CF dataset, detailed in the Appendix  B.1.2. By applying the SOTA locate-then-edit method PMET to layer [3-8], which follows (Li et al., 2024c), we present the percentage of cases where both pre-edited and post-edited models answer successfully in QA format or Cloze format under different edit batches.
Table 1 shows the results of the comparative experiments. We can see that performance on $D_{Pre}$ is significantly better than on $D_{Post}$, which aligns with our expectations. This is because reasoning the first hop knowledge in $D_{Pre}$ is similar to the single-hop process. After updating the knowledge in the earlier MLP layers, the model is likely to effectively use the newly edited knowledge. It can use the new implicit subject in the second hop to produce the final updated answer. When facing cases in $D_{post}$, PMET cannot get the correct final answer because it only modifies the earlier MLP layers, which is not enough for the model to correctly reason the second hop knowledge as it should be retrieved from later MLP layers.
4 IFMET: An Advanced Locate-then-Edit Method
Motivated by our findings on the distinctions between single-hop and multi-hop factual recall process, we introduce the Interpretability-Guided Furtherance Model Editing in a Transformer (IFMET). This method addresses the limitations identified in existing locate-then-edit approaches by modifying knowledge across both earlier and later MLP layers, enhancing the model’s ability to handle multi-hop reasoning. The IFMET method comprises two main steps: first, constructing a supplementary set of original edits to enrich the edit context, and second, performing editing based on multi-hop prompts derived from the original edit case and its supplementary set. This furtherance step approach ensures a thorough integration of new knowledge, significantly improving the model’s accuracy and robustness in multi-hop factual recall scenarios.
@@ 4 IFMET: An Advanced Locate-then-Edit Method
$s$ Spain
$T_{r}(s)$ The capital city of Spain is
$C$ (Manuel Almunia, cizizenship, Spain)
(Spain, capital, Madrid $\to$ Hartford)
$Q_{C}(s)$ What is the capital city of the country
where Manuel Almunia holds citizenship?
$e^{\prime}$ (Barcelona, country, Spain)
$s^{\prime}$ Barcelona
$C^{\prime}$ (Barcelona, country, Spain)
(Spain, capital, Madrid $\to$ Hartford)
$T_{C}(s^{\prime})$ The capital city of the country
@@
Table 2: An support case for an instance in the MQuAKE-CF dataset and the corresponding additional support cases are shown in the lower part.
Supplementary Set Construction. Note that for a given edit $e=\left(s,r,o\to o^{*}\right)$ (it can be extended to cases involving multiple edited facts), a locate-then-edit algorithm typically aims to identify and modify the knowledge-storing MLPs. Previous efforts have predominantly focused on the earlier MLP layers; however, our findings indicate that such an approach underperforms when the edited knowledge appears in second or subsequent hops during reasoning. Given that each edit traditionally targets single-hop knowledge, our experiments have demonstrated that using such edit prompts alone does not effectively update the later knowledge-storing MLPs. To address this issue, we construct a supplementary set for each edit, designed to facilitate the modification of deeper MLPs that provide knowledge in implicit fact recall steps.
In our supplementary set, we transform each edit into a multi-hop chain. For instance, for an edit $e=(s,r,o\to o^{*})$, we can create a supplementary fact $e_{\text{sup}}=(s^{\prime},r^{\prime},o^{\prime})$ where $o^{\prime}=s$, forming a two-hop fact recall chain $C=(s^{\prime},r^{\prime},o^{\prime})\oplus(s,r,o)$. This approach enables us to subsequently target and modify the latter MLPs that store the fact $(s,r,o)$, updating the information to $(s,r,o^{*})$. An illustrative example of this process is provided in Table 2.
Practically, we utilize WikiData333www.wikidata.org to construct the supplementary dataset. We start by extracting all subjects from the dataset’s edits and deduplicating them to form a set of subjects $S_{e}=\{s_{i}|i=1,\dots\}$. We then perform a WikiData SPARQL query [^4] https://query.wikidata.org/ to identify a set of triplets for each subject $s_{i}$: $Sup=\{(s^{\prime},r^{\prime},o^{\prime})|o^{\prime}=s_{i}\}$. To ensure the reliability of these facts, we filter out examples that cannot be correctly answered using the few-shot approach proposed by (Zhong et al., 2023). For construction details, please refer to the Appendix C.
Interpretability-Enhanced Furtherance Model Editing in a Transformer. Now we introduce the proposed IFMET framework. Each pre-edited knowledge has an additional multi-hop chain, assisted by the supplementary set. Based on the difference between the single and multi-top settings we discussed above, we have to locate and modify weights in both earlier and later layers in MLPs.

### v3

@@ 3.1 How Pre-edited Models Reason Fact Recall Tasks
For a multi-hop fact recall task, the model may employ multiple strategies to answer such tasks, including the formation of a single super-relation (Ju et al., 2024) $(s_{1},r_{mul},o_{n})$, where $r_{mul}=r_{1}\to\dots\to r_{n}$, or by segmenting the task into one explicit recall step followed by several implicit recall steps to answer step-by-step. Previous research (Hou et al., 2023) suggests that models typically engage in reasoning by considering each single-hop recall individually.
Based on this, we hypothesize that the model prioritizes deducing the implicit subject $s_{2}$ and subsequently recalls the final answer $o_{2}$ based on it. The following sections aim to verify this hypothesis by addressing the three questions: Q1: In the case of multi-hop reasoning, is the information related to $s_{2}$ accumulated before that of $o_{2}$? Q2: Does the accumulation of relevant information about $s_{2}$ causally influence the model’s reasoning for $o_{2}$? Q3: Which component facilitates the influence of factual recall process from $s_{2}$ to $o_{2}$?
@@
Is the information related to $s_{2}$ accumulated before that of $o_{2}$ We use LogitLens to examine the accumulation of information related to the implicit subject $s_{2}$ and the final answer $o_{2}$ in the two-hop scenario. The model’s predictions for $o_{2}$, are derived from the last token of the prompt, where crucial information about the resolved implicit subject $s_{2}$ should be propagated (Biran et al., 2024). Therefore, we focus on the hidden state $h_{l}$ at the $l$-th layer of the last token position, analyzing $\text{Info}(h_{l},s_{2})$ and $\text{Info}(h_{l},o_{2})$ as measures of the information related to $s_{2}$ and $o_{2}$ contained in $h_{l}$.
@@
The results, depicted in Figure 2(a), show that $\text{Info}(h_{l},s_{2})$ gradually reaches its peak during middle layers [15-17], while $\text{Info}(h_{l},o_{2})$ increases and peaks during later layers [21-24]. This pattern suggests that, in multi-hop tasks, the implicit subject $s_{2}$ is processed during the middle layers before reaching the final answer $o_{2}$. We conducted a similar experiment by giving $s_{2}$ explicitly in a single-hop prompt. The results, shown in Figure 2(b), indicate that there is no significant peak for the subject information before the final answer probability begins to accumulate, suggesting that in single-hop cases, the accumulation process of the final answer at the last token is not significantly correlated with the subject information.
@@
Does the accumulation of relevant information about $s_{2}$ causally influence the model’s reasoning for $o_{2}$? We propose an intervention experiment where we reduce the information content of $s_{2}$ at the last token position and observe changes in the output probability of the final answer in the last prediction layer.
Specifically, we replace the hidden state $h_{l}$ (in layer $\ell$ of the last token) with $h^{*}_{l}$, and the corresponding logits $s_{l}$ ($=W_{U}h_{l}$) changes to $s^{*}_{l}$ ($=W_{U}h^{*}_{l}$). $s^{*}_{l}$ is defined as:
where we minimize the logits corresponding to the tokens in $s_{2}$ without altering the logits of other tokens, aiming to diminish the effect of $s_{2}$. This setup allows us to describe the process through a causal intervention framework, where the target $\mathcal{H}$ is $h_{l}$, the intervention $\mathcal{I}_{h}$ and the effect $IE_{h}$ are defined as follows:
$\mathcal{I}_{h}:h^{*}_{l}=h_{l}+\operatorname*{arg\,min}_{\Delta h_{l}}\|W_{U}(h_{l}+\Delta h_{l})-s^{*}_{l}\|^{2},$
$IE_{h}=p_{L}[j]-p_{L}^{E}[j],\quad j\in o_{2},$ (2)
where $L$ is the last layer, $p_{L}[j]$ denotes the original output probability of $o_{2}$ in the $L$-th layer, and $p^{E}_{L}[j]$ is the probability after the intervention is applied. This approach illustrates how the hidden states and probabilities are expected to change when the logits are modified to $s^{*}$. For computational efficiency, we opt to approximate $h^{*}_{l}$ using a combination of least squares and minimum-norm methods (Lawson & Hanson, 1995) (further details are provided in Appendix C.1). For comparison, we also randomly select an irrelevant token $j\notin s_{2}\cup o_{2}$ to execute the intervention as the control group.
Figure 3(a) presents the outcomes of our intervention experiments across all layers, where a brighter color signifies a stronger intervention effect. We found a clear positive impact from intervening in layers [17-19] for the experimental group, in contrast to no significant effects observed in the control group across all layers. This suggests that, in the last token position, the information of $s_{2}$ encoded in the intermediate layers plays a crucial role in the probability accumulation process of $o_{2}$. We also do the same causal intervention experiments for single-hop fact recall, which suggest that the prediction of $o_{2}$ at the last token position is largely independent of $s_{2}$(see Appendix  C.2 for the results).
Takeaway 2 Unlike the mechanism of reasoning the knowledge in single-hop scenarios, in the reasoning process of the second-hop knowledge in two-hop scenarios, the accumulated implicit subject information has causal effects on the final answer.
Which component facilitates the influence of the factual recall process from $s_{2}$ to $o_{2}$? As previous studies claimed that single-hop tasks using subject information to retrieve knowledge from MLP layers (Meng et al., 2022a, b), we focus on MLP layers to answer how the implicit subject $s_{2}$ influences the prediction of the final answer $o_{2}$ in multi-hop scenario. We conducted a causal intervention experiment similar to the experiments above but focused specifically on the MLP component. Specifically, we aim to replace $m_{l}$ (the output hidden state of the last token in the $l$-th MLP layer) with $m^{*}_{l}$, where we have $s_{l}=W_{U}m_{l}$ and $s^{*}_{l}=W_{U}m^{*}_{l}$ with $s^{*}_{l}$ is same as in (1). The intervention $\mathcal{I}_{m}$ shares the same idea as in (3.1), except that $h_{l}$ is replaced with $m_{l}$. However, we redefine the intervention effect $IE_{m}$, which differs from the previous $IE_{h}$. In detail, we alternate the metric to the probability calculated from the output of MLP at the modified layer $l$ instead of the last layer $L$. In total, our causal intervention is formulated as
$\mathcal{I}_{m}:m^{*}_{l}=m_{l}+\operatorname*{arg\,min}_{\Delta m_{l}}\|W_{U}(m_{l}+\Delta m_{l})-s^{*}_{l}\|^{2},$
$IE_{m}=p_{l}[j]-p_{l}^{E}[j],\quad j\in o_{2}.$
@@
(b) Causal Intervention result of MLP hidden state in last token position.
@@
Figure 3: Causal Intervention Result: A brighter color signifies a stronger intervention effect. Note that negative effect values ($\leq$ 0) are clipped to 0 in both groups for better visualization. (a) is probability change $IE_{h}$ of intervention $\mathcal{I}_{h}$, (b) is probability change $IE_{m}$ of intervention $\mathcal{I}_{m}$.
@@
Figure 3(b) presents the outcomes of our intervention experiments across all layers with a similar control group as in the above. The clear positive impact from intervening in the intermediate layers [17-21] is demonstrated in the experimental group, in contrast to negligible effects observed in the control group across all layers. This suggests that the implicit subject $s_{2}$ at the last token position was used for retrieving the related information of $o_{2}$ from deeper MLP layers. Note that previous work (Meng et al., 2022a, 2023) has mentioned that explicit single-hop tasks primarily rely on the subject token position to retrieve information from shallow MLP layers.
Takeaway 3 In single-hop fact recall, relevant knowledge is retrieved through the token at the subject position, utilizing shallow MLP layers. In contrast, when the same knowledge serves as an implicit reasoning step in multi-hop, it is retrieved through the token at the last position, utilizing deeper MLP layers.
@@ 3.2 Why Existing locate-then-edit KE Methods Failed
Based on the findings above, we can explain the unsatisfactory performance of the existing locate-then-edit methods. The factual recall mechanism for the same knowledge differs when it serves as a single-hop reasoning step versus an implicit reasoning step in multi-hop reasoning. Consequently, for an editing instance $(s,r,o\to o^{*})$, using only the corresponding explicit single-hop prompt for editing in previous methods is insufficient as they only update the relevant knowledge in the shallow MLP layers but fail to propagate the changes to deeper layers, which are essential for multi-hop factual recall tasks.
We provide a concrete example in Table 1 for a better understanding. Given an editing instance $e$ and $T_{e}$. Existing methods modify the weights of shallow MLPs with $T_{e}$ to make it answer Hartford. $C_{pre}$ and $C_{post}$ represent the multi-hop factual recall chains obtained by $e$ as an explicit recall step and an implicit recall step, respectively. $T_{C_{pre}}$ and $T_{C{post}}$ represent the corresponding prompts. In this example, the query $T_{C_{pre}}$ should be answered correctly because the explicit fact $(Spain,capital,Hartford)$ can be recalled in shallow MLPs. However, the $T_{C_{post}}$ is still answered with Madrid because the knowledge $(Spain,capital,Madrid)$ stored in deeper MLPs does not change. To verify our claim above, we divide the two-hop factual recall tasks into two sets, $D_{Pre}$ and $D_{Post}$, depending on the position of the edited knowledge within the two-hop reasoning process. Specifically, for an edit instance $e=(s,r,o,o^{*})$, we have the following two sets:
$D_{Pre}=\{\left(s,r,o^{*}\right)\oplus\left(s_{2},r_{2},o_{2}\right)\},$
$D_{Post}=\{\left(s_{1},r_{1},o_{1}\right)\oplus\left(s,r,o^{*}\right)\}.$
The specific experimental details are in the appendix C.3. Table 2 presents the results of the comparative experiments. As shown, the multi-hop question accuracy performance of the existing SOTA locate-then-edit method PMET on $D_{\text{Pre}}$ is significantly better than on $D_{\text{Post}}$, which aligns with our expectations. This is because the reasoning in the explicit recall step is similar to the single-hop process. After updating the knowledge in the shallow MLP layers by single-hop edit prompt, the newly edited knowledge can be leveraged in $D_{pre}$. In contrast, for cases in $D_{\text{Post}}$, the model fails to produce the correct final answer because existing methods didn’t update the knowledge in deeper MLP layers.
@@
$T_{e}$ The capital city of Spain is
$C_{pre}$ e $\oplus$ (Hartford, mayor, Arulampalam)
$T_{C_{pre}}$ The mayor of the capital city
of Spain is
$C_{post}$ (Barcelona, country, Spain) $\oplus$ e
$T_{C_{post}}$ The capital city of the country
@@
Table 1: An example for an single-hop edit instance and its corresponding multi-hop prompt.
Table 2: Comparison of multi-hop Acc for $D_{Pre}$,$D_{Post}$.
Edit Method QA Format(%)    $\uparrow$ Cloze Format(%)    $\uparrow$
$D_{Pre}$ $D_{Post}$ $D_{Pre}$ $D_{Post}$
Base 50.62 41.72 20.31 18.63
PMET 64.29 2.93 43.37 4.60
Editor +CoT Average # Edits = # hops =
1-edit 2-edit 3-edit 4-edit 2-hop 3-hop 4-hop
Base 42.83 36.96 45.27 46.85 48.51 48.9 30.7 48.9
FT 1.9 4.2 0.7 0.3 0.0 3.7 1.4 0.5
MEND 11.5 16.0 11.0 7.3 4.4 13.9 11.3 9.5
ROME 18.1 23.8 20.9 9.0 2.6 33.8 9.1 11.4
MEMIT 12.3 20.5 9.8 5.5 2.6 22.5 6.0 8.4
PMET 17.04 22.63 16.74 11.19 7.84 26.65 12.76 11.7
IFMET (ours) 31.01 30.26 35.21 24.30 31.72 44.06 23.58 25.4
Table 3: Multi-hop accuracy comparison of different methods on the MQuAKE-3K dataset in a few-shot setting, showing the Base model’s performance on the unedited answer and the edited model’s performance on the edited answer.
4 IFMET
Motivated by our findings on the distinctions between single-hop and multi-hop factual recall processes, we introduce the Interpretability-Guided Furtherance Model Editing in a Transformer (IFMET). IFMET extends the existing locate-then-edit paradigm in two ways: first, constructing multi-hop edit prompts for each edit instance to expand the original single-hop edit prompt, and second, adding a furtherance editing step that applies multi-hop edit prompts to deeper MLPs. IFMET thoroughly integrates new knowledge across shallow and deeper MLP layers, significantly improving the model’s accuracy and robustness in multi-hop factual recall scenarios.
Multi-hop edit prompt construction. For a given edit $e=\left(s,r,o\to o^{*}\right)$, existing locate-then-edit methods provide only a single-hop edit prompt such as $T_{e}$ in Table 1. Through Section 3.2, we recognize that the main limitation of these methods lies in ignoring the difference between explicit single-hop and implicit multi-hop reasoning mechanisms. Single-hop edit prompts only modify knowledge in the shallow MLP layers, leading to poor performance in the $C_{post}$ format multi-hop factual recall. Therefore, we aim to construct $T_{C{post}}$-format multi-hop edit prompt for each edit instance, just like mentioned in Table 1. We first transform each edit instance into a two-hop fact recall chain $C=(s^{\prime},r^{\prime},o^{\prime})\oplus(s,r,o)$ where $o^{\prime}=s$, then transform this two-hop factual recall chain into an edit prompt by using the corresponding prompt templates. The key step is to identify the relevant preceding knowledge $(s^{\prime},r^{\prime},s)$ for $e$, which can be sourced from any valid knowledge base. Here, we provide two methods: one based on WikiData and the other using the model itself. Further discussion of these two construction methods is detailed in Appendix D.1.
IFMET. Now we introduce the proposed IFMET framework, providing a single-hop edit prompt and multi-hop edit prompt for each edit instance. Based on the difference between the single and multi-top reasoning mechanisms we discussed above, in the first edit stage, we use the single-hop edit prompt to edit shallow MLPs. In the second stage, we further use the multi-hop edit prompt to edit deeper MLPs.